OpenStudy (anonymous):

how do we find the fifth root of .0076 using log. Pls explain in full steps!!!!!!!!!

OpenStudy (anonymous):

you mean like with a calculator?

OpenStudy (anonymous):

old style method would be to say \[\log(\sqrt[5]{.0076})=\frac{1}{5}\log(.0076)\] but to use this you would have to know both the log of .0076 and also the antilog of your answer is this a history lesson?

OpenStudy (anonymous):

in other words if you knew the log of .0076 you would divide that number by 5, then take the antilog of the restult using a table in a very old dusty book that no one owns any more

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

\[\log_{.0076}X={1\over5} \]

OpenStudy (anonymous):